Session: Oct 2016 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (G.Maccaferri at Thu Oct 20 16:17:10 2016) Success:

Yebes (Ruben Fernández at Thu Oct 20 16:26:22 2016) Success:
No known problem.

Tianma65 (BoXia at Fri Oct 21 00:34:01 2016) Success:

Irbene (Vladislavs Bezrukovs at Fri Oct 21 13:30:31 2016) Minor failures: RECO FS
Strange offset in time in disk2file: 14 - 16 seconds offset. FS crashed, 9 - 10 scans lost.

Westerbork (Geert Kuper at Fri Oct 21 14:43:33 2016) Success:

Onsala (Jun Yang at Mon Oct 24 07:08:05 2016) Success:
Fringes were clearly seen during the ftp-fringe tests.

Noto (P.Cassaro at Tue Oct 25 22:41:14 2016) Success:

Effelsberg (Uwe at Wed Nov 02 11:42:16 2016) Minor failures:
Due to a power failure an hour before the start, that caused a failure of the receiver cooling, observations started somewhat later and ended 30min earlier. An amplifier in LCP seems to be affected as well, Tsys is 10x of the RCP value. Therefore, the LCP SNR in the fringe test was about 3x lower than usual.